# Neural Deconstruction Search for Vehicle Routing Problems

### v1

@@ Abstract
Autoregressive construction approaches generate solutions to vehicle routing problems in a step-by-step fashion, leading to high-quality solutions that are nearing the performance achieved by handcrafted, operations research techniques. In this work, we challenge the conventional paradigm of sequential solution construction and introduce an iterative search framework where solutions are instead deconstructed by a neural policy. Throughout the search, the neural policy collaborates with a simple greedy insertion algorithm to rebuild the deconstructed solutions. Our approach surpasses the performance of state-of-the-art operations research methods across three challenging vehicle routing problems of various problem sizes.
@@ 1 Introduction
NDS is an iterative search method designed to enhance a given solution through a two-phase process involving deconstruction and reconstruction along the lines of large neighborhood search (LNS) (Shaw, 1998) and ruin-and-recreate (Schrimpf et al., 2000) paradigms. The deconstruction phase employs a deep neural network (DNN) to determine the customers to be removed from the tours of the current solution. This is achieved through a sequential decision-making process, in which nodes are removed one at a time based on the network’s guidance. The reconstruction phase utilizes a straightforward greedy insertion algorithm, which inserts customers in the order given by the neural network at the locally optimal positions. The core concept of NDS is shown in Figure 1. Note that NDS is trained using reinforcement learning, which makes it adaptable to problems for which no reference solutions are available for training.
@@
The overall concept of modifying a solution by first removing some solution components and then conducting a rebuilding step has been successfully used in various vehicle routing problem methods. Non-learning based methods that use this concept include the rip-up and reroute method from Dees & Smith (1981), LNS from Shaw (1998), and the ruin and recreate method from Schrimpf et al. (2000). Learning-based methods have also harnessed this paradigm. The local rewriting method from Chen & Tian (2019), neural large neighborhood search from Hottung & Tierney (2020), and the random reconstruction technique introduced in Luo et al. (2023) employ a DNN during the reconstruction phase. The approaches from Li et al. (2021) and Falkner & Schmidt-Thieme (2023) both generate different subproblems for a given solution and then use a DNN to choose which subproblem should be considered in the reconstruction phase.
@@
NDS has been designed with the goal of achieving a fast search procedure without sacrificing the high-quality search guidance of a DNN. For medium-sized CVRP instances with 500 customers, state-of-the-art OR approaches such as SISRs (Christiaens & Vanden Berghe, 2020) can examine upwards of 270k solutions per second, however neural combinatorial optimization approaches, like POMO (Kwon et al., 2020), can only observe around 10k per second. In contrast, NDS can process 120k solutions per second, significantly more than existing neural construction techniques. When combined with a powerful deconstruction DNN, NDS is able to outperform state-of-the-art OR approaches like SISRs and HGS in similar wall-clock time.
Figure 1: Improving a solution via neural deconstruction.
@@
We evaluate NDS on several challenging problems, including the CVRP, the vehicle routing problem with time windows (VRPTW), and the price-collecting vehicle routing problem (PCVRP). NDS demonstrates substantial performance gains compared to existing learned construction methods and surpasses state-of-the-art OR methods across various routing problems of different sizes. To the best of our knowledge, NDS is the first learning-based approach that achieves this milestone.
@@ 2 Literature Review
Various techniques have been proposed to enhance performance in neural combinatorial optimization. For instance, Hottung et al. (2022) introduce efficient active search, which updates a subset of parameters during inference. Choo et al. (2022) propose SGBS, combining Monte Carlo tree search with beam search to guide the search process more effectively. Additionally, Drakulic et al. (2023) and Luo et al. (2023) focus on improving out-of-distribution generalization by re-encoding the remaining subproblem after each construction step. To enhance solution diversity during sampling, Grinsztajn et al. (2022) and Hottung et al. (2024) explore approaches that learn a set of policies, rather than a single policy.
Instead of constructing solutions autoregressively, some approaches predict heat maps that highlight promising solution components (e.g., arcs in a graph), which are then used in post-hoc searches to construct solutions (Joshi et al., 2019; Fu et al., 2021; Kool et al., 2022b; Min et al., 2023). Other approaches focus on more complex variants of routing problems, such as the VRPTW (Falkner & Schmidt-Thieme, 2020; Kool et al., 2022a; Berto et al., 2024a; b).
@@
Improvement methods focus on iteratively refining a given starting solution. In addition to the ruin-and-recreate approaches discussed in the introduction, several other methods aim to enhance solution quality through iterative adjustments. For instance, Ma et al. (2021) propose learning to iteratively improve solutions by performing local modifications. Similarly, several works have guided the $k$-opt heuristic for vehicle routing problems (Wu et al., 2019; da Costa et al., 2020), although they are constrained by a fixed, small $k$. More recently, Ma et al. (2023) introduced a method capable of handling any $k$. Furthermore, Ye et al. (2024a) and Kim et al. (2024) integrate learning-based approaches with ant colony optimization to allow for a more extensive search phase. Additionally, several divide-and-conquer methods have been developed to address large-scale routing problems (Kim et al., 2021; Li et al., 2021; Ye et al., 2024b).
@@ 3.1 Deconstruction Policy
For solution deconstruction, a neural policy is employed to sequentially select customers for removal from a given solution. We model this selection process as a Markov decision process. Let $s$ be a feasible solution to a vehicle routing problem (VRP) instance $l$, which involves customers $c_{1},\dots,c_{N}$. A policy network $\pi_{\theta}$, parameterized by $\theta$, is used to select $M$ customers for removal. At each step $m\in\{1,\dots,M\}$, an action $a_{m}\in\{1,\dots,N\}$ is chosen according to the probability distribution $\pi_{\theta}(a_{m}\mid l,s,v,a_{1:m-1})$, where $a_{m}$ corresponds to selecting customer $c_{a_{m}}$, $l$ is the instance, $s$ is the solution, $v$ is a random seed, and $a_{1:m-1}$ are the previous actions. We condition the policy on a random seed $v$ to encourage more diverse rollouts as explained in Hottung et al. (2024). Each seed is a randomly generated binary vectors of dimension $d_{v}$ (we set $d_{v}=10$ in all experiments). Finally, after all $M$ customers are selected the reward can be computed as discussed in the following sections.
@@ 3.2 Training
7:              $s\leftarrow\textsc{ImprovementStep}(s,\pi_{\theta})$triangleright Improve solution using procedure shown in Figure 2
@@
10:              $\{\tau_{1},\tau_{2},\dots,\tau_{K}\}\leftarrow\textsc{RolloutPolicy}(\pi_{\theta},l,s,K)$
@@
13:              $r_{k}\leftarrow\max(\textsc{Obj}(s)-\textsc{Obj}(s^{\prime}_{k}),0)\qquad\forall k\in\{1,\dots,K\}$triangleright Calculate reward
14:              $b\leftarrow\frac{1}{K}\sum_{k=1}^{K}r_{k}$triangleright Calculate baseline
@@
16:              $g_{i}\leftarrow(r_{k^{*}}-b)\nabla_{\theta}\log\pi_{\theta}(\tau_{k^{*}}|l,s,v_{k^{*}})$triangleright Calculate gradients
17:              $s\leftarrow s^{\prime}_{k^{*}}$triangleright Update $s$ with best found solution
@@
19:         $\theta\leftarrow\theta+\alpha\sum_{i=1}^{I}g_{i}$triangleright Optimizer step with accumulated gradients
@@
At the start of each iteration of the training loop, a new instance $l$ and its corresponding solution $s$ are generated. The instance is sampled from the same distribution as the test instances. In line 5, an initial solution is constructed using a simple procedure: for an instance with $N$ customers, we generate $N$ tours, each containing one customer. In lines 6-8, this initial solution is iteratively improved through $J$ improvement steps of the NDS search procedure, which are detailed in Section 3.4. By refining $s$ with NDS’s main search component before the training rollouts, we ensure that the training focuses on improving non-trivial solutions.
@@
The greedy insertion procedure reintegrates the customers removed by the policy, inserting them one by one into either existing or new tours. Specifically, if $M$ customers have been removed, the procedure performs $M$ iterations, where in each iteration, a single customer $c_{a_{m}}$ is inserted. At each iteration $m$, the cost of inserting customer $c_{a_{m}}$ at every feasible position in the current tours is evaluated. Throughout this process, various constraints, such as vehicle capacity limits, are taken into account. If at least one feasible insertion point is found within an existing tour, the customer $c_{a_{m}}$ is placed at the position that incurs the least additional cost. If no feasible insertion is available, a new tour is created for customer $c_{a_{m}}$.
The order in which removed customers are reinserted significantly impacts the overall performance. We reinsert customers either in the order determined by the neural network or at random. Allowing the network to control the reinsertion order gives it control over the reconstruction process, enabling it to find ordering strategies that lead to better reconstructed solutions. If customers are ordered at random, a deconstructed solution should be reconstructed multiple times using different insertion orders. This can provide more stable learning signal during training.
@@ 3.3.1 Encoder
The encoder processes the input features $x_{i}$ for each of the $N+1$ nodes, where $x_{0}$ corresponds to the depot’s features, alongside the current solution $s$ that needs to be encoded. Initially, each input vector $x_{i}$ is mapped to a 128-dimensional node embedding $h_{i}$ through a linear transformation. The embeddings $h_{0},\dots,h_{N}$ are sequentially processed through several layers. First, two attention layers encode static instance information. Next, a message passing layer allows information exchange between consecutive nodes in the solution. This is followed by a tour embedding layer, which computes embeddings for each tour within the solution. Finally, two additional attention layers refine the representations. The attention mechanisms employed are consistent with those used in prior work (e.g., Kwon et al. (2020)), and detailed descriptions are omitted here for brevity.
@@ 3.3.2 Decoder
Given the node embeddings generated by the encoder, the decoder is responsible for sequentially selecting customers for removal. The overall architecture of our decoder is identical to that of Hottung et al. (2024), which utilizes a multi-head attention mechanism (Vaswani et al., 2017) followed by a pointer mechanism (Vinyals et al., 2015). This architecture has been widely used in many routing problems methods (Kool et al., 2019; Kwon et al., 2020).
Our approach differs from previous works in that we account for the already selected customers at each decision step. This contrasts with construction-based methods, where each decision is independent of prior selections. To address this, we integrate a gated recurrent unit (GRU) (Cho, 2014), which is used to compute the query for the multi-head attention mechanism. At each decision step, the GRU takes the embedding of the previously selected customer as input, updating its internal state to incorporate past decisions.

### v2

@@ Abstract
Autoregressive construction approaches generate solutions to vehicle routing problems in a step-by-step fashion, leading to high-quality solutions that are nearing the performance achieved by handcrafted operations research techniques. In this work, we challenge the conventional paradigm of sequential solution construction and introduce an iterative search framework where solutions are instead deconstructed by a neural policy. Throughout the search, the neural policy collaborates with a simple greedy insertion algorithm to rebuild the deconstructed solutions. Our approach matches or surpasses the performance of state-of-the-art operations research methods across three challenging vehicle routing problems of various problem sizes.
@@ 1 Introduction
NDS is an iterative search method designed to enhance a given solution through a two-phase process involving deconstruction and reconstruction along the lines of large neighborhood search (LNS) (Shaw, 1998) and ruin-and-recreate (Schrimpf et al., 2000) paradigms. The deconstruction phase employs a deep neural network (DNN) to determine the customers to be removed from the tours of the current solution. This is achieved through a sequential decision-making process, in which nodes are removed one at a time based on the network’s guidance. The reconstruction phase utilizes a straightforward greedy insertion algorithm, which inserts customers in the order given by the neural network at the locally optimal positions. Note that NDS is trained using reinforcement learning, which makes it adaptable to problems for which no reference solutions are available for training.
@@
The overall concept of modifying a solution by first removing some solution components and then conducting a rebuilding step has been successfully used in various vehicle routing problem methods. Non-learning-based methods that use this concept include the rip-up and reroute method from Dees & Smith (1981), LNS from Shaw (1998), and the ruin and recreate method from Schrimpf et al. (2000). Learning-based methods have also harnessed this paradigm. The local rewriting method from Chen & Tian (2019), neural large neighborhood search from Hottung & Tierney (2020), and the random reconstruction technique introduced in Luo et al. (2023) employ a DNN during the reconstruction phase. The approaches from Li et al. (2021) and Falkner & Schmidt-Thieme (2023) both generate different subproblems for a given solution and then use a DNN to choose which subproblem should be considered in the reconstruction phase.
@@
NDS has been designed with the goal of achieving a fast search procedure without sacrificing the high-quality search guidance of a DNN. For medium-sized CVRP instances with 500 customers, state-of-the-art OR approaches such as SISRs (Christiaens & Vanden Berghe, 2020) can examine upwards of 270k solutions per second, however neural combinatorial optimization approaches, like POMO (Kwon et al., 2020), can only create around 10k per second. In contrast, NDS can process 120k solutions per second, significantly more than existing neural construction techniques. When combined with a powerful deconstruction DNN, NDS is able to outperform state-of-the-art OR approaches like SISRs and HGS in similar wall-clock time.
@@
We evaluate NDS on several challenging problems, including the CVRP, the vehicle routing problem with time windows (VRPTW), and the prize-collecting vehicle routing problem (PCVRP). NDS demonstrates substantial performance gains compared to existing learned construction methods and matches or surpasses state-of-the-art OR methods across various routing problems of different sizes. To the best of our knowledge, NDS is the first learning-based approach that achieves this milestone.
@@ 2 Literature Review
Various techniques have been proposed to enhance performance in neural combinatorial optimization. For instance, Hottung et al. (2022) introduce efficient active search, which updates a subset of model parameters during inference. Choo et al. (2022) propose SGBS, combining Monte Carlo tree search with beam search to guide the search process more effectively. Additionally, Drakulic et al. (2023) and Luo et al. (2023) focus on improving out-of-distribution generalization by re-encoding the remaining subproblem after each construction step. To enhance solution diversity during sampling, Grinsztajn et al. (2023) and Hottung et al. (2025) explore approaches that learn a set of policies, rather than a single policy.
Instead of constructing solutions autoregressively, some approaches predict heat maps that highlight promising solution components (e.g., arcs in a graph), which are then used in post-hoc searches to construct solutions (Joshi et al., 2019; Fu et al., 2021; Kool et al., 2022b; Min et al., 2023). Other approaches focus on more complex variants of routing problems, such as the VRPTW (Falkner & Schmidt-Thieme, 2020; Kool et al., 2022a; Berto et al., 2024b; c), or the min-max heterogeneous CVRP (Berto et al., 2024a).
@@
Improvement methods focus on iteratively refining a given starting solution. In addition to the ruin-and-recreate approaches discussed in the introduction, several other methods aim to enhance solution quality through iterative adjustments. For instance, Ma et al. (2021) propose learning to iteratively improve solutions by performing local modifications. Similarly, several works have guided the $k$-opt heuristic for vehicle routing problems (Wu et al., 2019; da Costa et al., 2020), although they are constrained by a fixed, small $k$. More recently, Ma et al. (2023) introduced a method capable of handling any $k$. Furthermore, Ye et al. (2024a) and Kim et al. (2024) integrate learning-based approaches with ant colony optimization to allow for a more extensive search phase. Additionally, several divide-and-conquer methods have been developed to address large-scale routing problems (Kim et al., 2021; Li et al., 2021; Ye et al., 2024b; Zheng et al., 2024).
3 Vehicle Routing Problems
Vehicle routing problems (VRPs) represent a broad class of combinatorial optimization problems that are fundamental in logistics and transportation. These problems involve determining optimal routes for a fleet of vehicles to serve a set of customers while satisfying specific constraints. The standard VRP generalizes the well-known TSP and is also NP-hard. In this paper, we consider three key VRP variants: the CVRP, the VRPTW, and the PCVRP. Each variant introduces additional constraints and objectives, making them suitable for different real-world applications. We briefly introduce all three variants below, while a more detailed discussion can be found in Toth & Vigo (2014).
Capacitated Vehicle Routing Problem
The CVRP is one of the most studied VRP variants, where each customer has a specific demand that must be met by a fleet of vehicles with limited capacity. Formally, the problem is defined on a complete graph where a depot and a set of $N$ customers, denoted as $c_{1},\dots,c_{N}$, are represented as nodes with associated coordinates in a two-dimensional Euclidean space. Each customer $c_{i}$ has a demand $q_{i}$, and each vehicle has a maximum capacity $Q$. The objective is to determine a set of routes that collectively serve all customers while minimizing the total travel cost. We calculate the travel costs as the sum of Euclidean distances between visited nodes. Each route must start and end at the depot, and the sum of customer demands on any route must not exceed $Q$. An instance of the problem is denoted by $l$, which encapsulates the locations of the depot and customers, their demands, and the vehicle capacity. A solution $s$ for an instance $l$ consists of a set of routes, where each route defines the sequence of customer visits assigned to one vehicle.
Vehicle Routing Problem with Time Windows
The VRPTW extends the CVRP by incorporating time constraints on customer deliveries. Each customer $c_{i}$ is assigned a time window $a_{i},b_{i}]$, specifying the earliest and latest allowable delivery times. A vehicle may arrive early but must wait until the time window opens. Additionally, each customer has a service time, representing the duration required to complete the delivery before the vehicle can proceed. The objective is to minimize total travel time while ensuring that all deliveries occur within the specified time windows. In this paper, the travel time between two nodes is identical to the Euclidean distance between those nodes. Like the CVRP, all routes start and end at the central depot, and vehicle capacity constraints must be respected.
Prize-Collecting Vehicle Routing Problem
The PCVRP is a variant of the VRP in which not all customers need to be visited. Unlike the CVRP and VRPTW, where all customers must be served, the PCVRP assigns a prize to each customer, and the objective is to maximize the difference between the total prize collected and the travel costs. Vehicles still start and end at a central depot and must respect capacity constraints. This variant is particularly useful in scenarios where serving every customer is not mandatory but must be balanced against operational costs.
@@ 4 Neural Deconstruction Search
NDS is an improvement method for vehicle routing problems. During the search, new candidate solution are generated by applying a neural deconstruction policy followed by a greedy reconstruction algorithm to the current solution. An example for generating a new candidate solution for the CVRP is shown in Figure 1. First, the static instance information $l$ and the current solution $s$ are given to a neural policy which then selects customers for removal from the tours of $s$ in a sequential decision making process. The selected customers are then removed from the tours of $s$, resulting in a (partial) solution in which the selected customers are not part of any tour and are thus unvisited. Finally, the selected customers are reinserted into tours using a simple sequential greedy insertion algorithm that inserts customers in the order defined by the policy.
Figure 1: Improving a solution via neural deconstruction.
Note that the number of decision steps (and hence the number of model calls) depends solely on the number of customers that are removed. Since we keep that number fixed across different instances sizes, NDS is able to generate candidate solutions much faster than construction approaches that build solutions from scratch. Furthermore, NDS is built for efficient GPU and CPU interaction by using batched rollouts.
One key requirement for the learned policies is that they contain a sufficient degree of randomness during the deconstruction process. During the search, a good solution is often deconstructed hundreds of times before an improving solution is found. It is hence vital that the learned policy is able to produce multiple diverse deconstruction instructions for a single solution. We encourage diverse output generation by providing the policy with an additional seed input $v$ as proposed in Hottung et al. (2025). During training, the model learns to condition its output on $v$, resulting in the selection of different customer sets for different seed values at test time.
@@ 4.1 Deconstruction Policy
For solution deconstruction, a neural policy is employed to sequentially select customers for removal from a given solution. We model this selection process as a Markov decision process. Given a solution $s$ for a VRP instance $l$, a policy network $\pi_{\theta}$, parameterized by $\theta$, is used to select $M$ customers for removal. At each step $m\in\{1,\dots,M\}$, an action $a_{m}\in\{1,\dots,N\}$ is chosen according to the probability distribution $\pi_{\theta}(a_{m}\mid l,s,v,a_{1:m-1})$, where $a_{m}$ corresponds to selecting customer $c_{a_{m}}$, $v$ is a random seed, and $a_{1:m-1}$ are the previous actions. We condition the policy on a random seed $v$ to encourage more diverse rollouts as explained in Hottung et al. (2025). Each seed is a randomly generated binary vectors of dimension $d_{v}$ (we set $d_{v}=10$ in all experiments). Finally, after all $M$ customers are selected the reward can be computed as discussed in the following sections.
@@ 4.2 Training
7:     $s\leftarrow\textsc{ImprovementStep}(s,\pi_{\theta})$ $\triangleright$ Improve solution using procedure shown in Figure 2
@@
10:     $\{\tau_{1},\tau_{2},\dots,\tau_{K}\}\leftarrow\textsc{RolloutPolicy}(\pi_{\theta},l,s,K)$ $\triangleright$ Sample $K$ rollouts
@@
13:     $r_{k}\leftarrow\max(\textsc{Obj}(s)-\textsc{Obj}(s^{\prime}_{k}),0)\qquad\forall k\in\{1,\dots,K\}$ $\triangleright$ Calculate reward
14:     $b\leftarrow\frac{1}{K}\sum_{k=1}^{K}r_{k}$ $\triangleright$ Calculate baseline
@@
16:     $g_{i}\leftarrow(r_{k^{*}}-b)\nabla_{\theta}\log\pi_{\theta}(\tau_{k^{*}}|l,s,v_{k^{*}})$ $\triangleright$ Calculate gradients
17:     $s\leftarrow s^{\prime}_{k^{*}}$ $\triangleright$ Update $s$ with best found solution
@@
19:   $\theta\leftarrow\theta+\alpha\sum_{i=1}^{I}g_{i}$ $\triangleright$ Optimizer step with accumulated gradients
@@
At the start of each iteration of the training loop, a new instance $l$ and its corresponding solution $s$ are generated. The instance is sampled from the same distribution as the test instances. In line 5, an initial solution is constructed using a simple procedure: for an instance with $N$ customers, we generate $N$ tours, each containing one customer. In lines 6-8, this initial solution is iteratively improved through $J$ improvement steps of the NDS search procedure. This search procedure is detailed in Section 4.4. By improving $s$ before the training rollouts, we ensure that the training focuses on non-trivial solutions.
@@
The greedy insertion procedure reintegrates the customers removed by the policy, inserting them one by one into either existing or new tours. Specifically, if $M$ customers have been removed, the procedure performs $M$ iterations, where in each iteration, a single customer $c_{a_{m}}$ is inserted. At each iteration $m$, the cost of inserting customer $c_{a_{m}}$ at every feasible position in all tours is evaluated. Throughout this process, various constraints, such as vehicle capacity limits, are taken into account. If at least one feasible insertion point is found within any of the existing tours, the customer $c_{a_{m}}$ is placed at the position that incurs the least additional cost. If no feasible insertion is available, a new tour is created for customer $c_{a_{m}}$.
The order in which removed customers are reinserted significantly impacts the overall performance. We reinsert customers either in the order determined by the neural network or at random. Allowing the network to control the reinsertion order gives it control over the reconstruction process, enabling it to find ordering strategies that lead to better reconstructed solutions. If customers are ordered at random, a deconstructed solution should be reconstructed multiple times using different insertion orders. This can provide a more stable learning signal during training.
@@ 4.3.1 Encoder
The encoder processes the static node features $x_{i}$ for each of the $N+1$ nodes in an instance $l$, where $x_{0}$ corresponds to the depot and $x_{1},\dots,x_{N}$ correspond to customer nodes. Additionally, it incorporates the current solution $s$. For the CVRP, the depot feature vector $x_{0}$ includes the depot coordinates, while the customer feature vectors $x_{i}$ ($i\geq 1$) include the customer coordinates and demands $q_{i}$. For the VRPTW, the customer feature vectors are further extended to include the earliest $a_{i}$ and latest arrival times $b_{i}$, as well as the service duration. For the PCVRP, the CVRP customer feature set is additionally extended to include customer prizes.
Initially, each input vector $x_{i}$ is mapped to a 128-dimensional node embedding $h_{i}$ via a linear transformation, using distinct parameters for the depot and customer nodes. The embeddings $h_{0},\dots,h_{N}$ are then processed through several layers. First, two attention layers encode static instance information. Next, the current solution $s$ is integrated via a message passing layer, followed by a tour encoding layer. The message passing layer facilitates information exchange between consecutive nodes in the solution, while the tour encoding layer computes embeddings for each tour. These two layers are explained in more detail below. Finally, two additional attention layers refine the representations. The attention mechanisms employed are consistent with those used in prior work (e.g., Kwon et al. (2020)), and detailed descriptions are omitted here for brevity.
@@ 4.3.2 Decoder
Given the node embeddings generated by the encoder, the decoder is responsible for sequentially selecting customers for removal. The overall architecture of our decoder is identical to that of Hottung et al. (2025), which utilizes a multi-head attention mechanism (Vaswani et al., 2017) followed by a pointer mechanism (Vinyals et al., 2015). This architecture has been widely used in many routing problem methods (Kool et al., 2019; Kwon et al., 2020).
Our approach differs from previous works in that we account for the already selected customers at each decision step. This contrasts with construction-based methods, where each decision is independent of prior selections. To achieve this, we integrate a gated recurrent unit (GRU) (Cho, 2014), which is used to compute the query for the multi-head attention mechanism. At each decision step, the GRU takes the embedding of the previously selected customer as input, updating its internal state to incorporate past decisions.